OpenStudy (anonymous):

another way to write y=3x+1?

OpenStudy (agent0smith):

Subtract 1 from both sides.

OpenStudy (anonymous):

Or there is standard form which would be -3x + y = 1 (you subtract 3x from both sides)

OpenStudy (anonymous):

thank you both!

OpenStudy (agent0smith):

Standard form would be 3x - y = -1 since usually the x coefficient is meant to be positive.